OpenStudy (anonymous):

GIVING MEDALS. simplify sqrt(1-cos theta)(1+cos theta)

OpenStudy (anonymous):

\[\sqrt{(1-\cos\theta)(1+\cos\theta)}~~?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Multiply. You have the following situation: \[(a-b)(a+b)=a^2-b^2\]

OpenStudy (anonymous):

okay whats next @SithsAndGiggles

OpenStudy (anonymous):

@mathstudent55 @RaphaelFilgueiras @Whitemonsterbunny17

OpenStudy (anonymous):

What's \((1-\cos\theta)(1+\cos\theta)\) equivalent to, using the difference of squares formula above?

OpenStudy (anonymous):

im not sure , answer choices

OpenStudy (anonymous):

\[(a-b)(a+b)=a^2-b^2\] In this case \(a=1\) and \(b=\cos\theta\).

OpenStudy (anonymous):

ookokay so which one

OpenStudy (anonymous):

can u help with this @SithsAndGiggles @ganeshie8

OpenStudy (anonymous):

do u know @ganeshie8